4. You established a straddle strategy for Tesla aiming for expiration date of this June. Tesla June 250 call contract has a premium of $18 and Tesla June 250 put contract has a premium of $17. Current stock price of Tesla is at 252. If Tesla’s stock price decreased to $200 by June, how much profit or loss you will realize on the investment? At what price range by expiration date will you lose money on your straddle strategy?
Answer #1

1.
=MAX(200-250,0)+MAX(250-200,0)-18-17=15

2.
Lower end=250-35=215
Higher end=250+35=285